1. A 50 KVA transformer has 400 turns on the primary and 50 turns on the secondary winding. The primary is connected to 2000 V, 50 Hz supply. Find
1) the full load primary and secondary currents
2) the secondary emf.
3) maximum flux in the core
1
Expert's answer
2020-11-09T14:18:25-0500

Following are given datas,

Transformer kVA"= 50kVA"

Primary turns"(N1) = 400"

Secondary turns"(N2) =50"

Primary input voltage "(V1) = 2000V"

Frequency"(f) =50Hz"


2) Secondary voltage :


"\frac{V2}{V1}=\frac{N2}{N1}"


"\implies V2=\frac{N2}{N1}\times V1
= \frac{50 }{400}\times2000= 250V"


1) Primary current "(I1) = \frac{kVA}{V1}=\frac{50000}{2000}=25A"


Secondary current"(I2) =\frac{kVA}{V2}= \frac{50000}{250}= 200A"


3) We have emf equation for transformer as follows,


"V1=4.44\phi f N1, Where \space \phi \space is \space Maximum flux"


There fore", \phi = \frac{V1}{4.44 f N1}= \frac{2000}{4.44\times 50\times400}= 0.0225 \space weber"